40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 7 22:58:08 2010, nancy, Update, IOO, MC alignment values. 
    Reply  Wed Jul 7 23:11:08 2010, Koji, Update, IOO, MC alignment values. 
       Reply  Thu Jul 8 20:32:22 2010, nancy, Update, IOO, MC alignment values. 
Message ID: 3183     Entry time: Thu Jul 8 20:32:22 2010     In reply to: 3175
Author: nancy 
Type: Update 
Category: IOO 
Subject: MC alignment values. 

I and Koji were trying to lock the mode cleaner for measuring the beam power at MC2 end. That is when we obtained the trans and refl values.

The beam characteristics at the MC2 were measured so that we could now use a dummy beam of similar power to test and characterize the QPD we are about to install at the MC2 end. This QPD wil provide two more signals in pitch and yaw, and hence complete 6 signals for 6 rotatioanl dof of the cavity. (4 are coming from WFS).

Once the QPD is characterised, it can be used to see the spot position at MC2. This is related to the mirror angles.

The width measurements were done using a beam scan. the beam scan was properly adjusted so that the maxima of the intensity of the sopt was at its center.

We also fitted gaussian curve to the beam profile, and it was a substantially good fit.

 

The whole idea is that I am trying  to look how the Wavefront sensors respond to the perturbations in the mirror angles. Once this is known, we should be able to control the mirror-movements.

the starting point would be to do just the DC measurements (which I did today). For proper analysis, AC measurements are obviously required.(will be done later).

The matrices so calculated can be inverted, and if found enough singular, the method can be used to control.

The first shot is to see teh dependency of teh error signals only on MC1 and MC3, and see if that is kind of enough to control these two mirrors.

If this works, the QPD signals could be used to control MC2 movements.

Quote:

Hmm. I expect that you will put more details of the work tomorrow.
i.e. motivation, method, result (the previous entry is only this),
and some discussiona with how to do next.

Quote:

Nancy and Koji:

This is what I and Koji measured after aligning the MC in the afternoon.

MC_Trans 4.595 (avg)

MC_Refl 0.203 (avg)

MC2_trans :

power = 1.34mW

13.5% width : x=6747.8 +- 20.7 um  , y = 6699.4+- 20.7 um

 

 

 

ELOG V3.1.3-